OpenStudy (anonymous):

need help; Find the equation of the line given a slope of 2/3 and a point (5, -4) ? so solve for y=mx+b in which I need b (y-intercept)

OpenStudy (tyteen4a03):

In this case, don't use the slope-intercept formula. Use the point slope form: \(y - y_1 = m(x - x_1)\)

OpenStudy (anonymous):

to find b?

OpenStudy (tyteen4a03):

@bpaquin007 No. Don't use the slope-intercept form at all. Use the point-slope form, then if required rebuild the equation to slope-intercept form.

OpenStudy (texaschic101):

I see no reason why you can't use y = mx + b and solve for b. Then write your equation of the line.

OpenStudy (anonymous):

whast the awnser then? I keep getting 14/3 = b

OpenStudy (anonymous):

its supposed to be y=2/3x = 22/3

OpenStudy (anonymous):

but I don't know how to get there

OpenStudy (texaschic101):

y = mx + b -4 = 2/3(5) + b -4 = 10/3 + b -4 - 10/3 = b -12/3 - 10/3 = b -22/3 = b your equation is : y = 2/3x - 22/3

OpenStudy (texaschic101):

are you still confused ?

OpenStudy (anonymous):

kind of

OpenStudy (anonymous):

can you explain it but maybe in words

OpenStudy (texaschic101):

y = mx + b slope(m) = 2/3 (5,-4) x = 5 and y = -4 sub your stuff into the formula y = mx + b -4 = 2/3(5) + b you understanding so far ?

OpenStudy (anonymous):

yep

OpenStudy (tyteen4a03):

@texaschic101 Why should you solve for b when you can plug in directly with another form of the same equation, then rebuild your way back to the original form of equation you want?

OpenStudy (tyteen4a03):

y + 4 = 2/3(x - 5) y = 2/3x - 10/3 + 4 y = 2/3x + 2/3 You can't get easier than that.

OpenStudy (tyteen4a03):

Uh: y = 2/3x - 10/3 - 4 y = 2/3x - 22/3 Whatever is wrong with my brain...

OpenStudy (texaschic101):

-4 = 2/3(5) + b -- distribute through the parenthesis -4 = 10/3 + b -- subtract 10/3 from both sides -4 - 10/3 = 10/3 - 10/3 + b as you can see, the 10/3's cancel out on the right side now we have : -4 - 10/3 = b -4 is the same as -4/1...we are dealing with fractions, so we need a common denominator, which is 3. Thus turning -4/1 into -12/3. Now we have : -12/3 - 10/3 = b add the two fractions and keep the negative sign -22/3 = b now put it into your equation.....remember m = 2/3 y = 2/3x - 22/3 does that help ?

OpenStudy (anonymous):

you there texas?

OpenStudy (texaschic101):

yep...I am here....I tried to explain in great detail...do you understand ?

OpenStudy (tyteen4a03):

@texaschic101 You are overcomplicating the solution. @bpaquin007 See my solution, which happens in 3 lines.

OpenStudy (anonymous):

THank you so much texas!!! I got it:)

OpenStudy (texaschic101):

The reason my solution is taking so long is because I am including every step

OpenStudy (texaschic101):

you can solve this by using y = mx + b or y - y1 = m(x - x1)...whichever way is easier for you to do.

OpenStudy (anonymous):

texas I have another question for you if your still here

OpenStudy (texaschic101):

lets see it :)

OpenStudy (anonymous):

find the equation of the line given these two points: (-2,-1) and (9,-3)

OpenStudy (texaschic101):

do you know how to find the slope ?

OpenStudy (anonymous):

yeah the slope is -2/11, you could skip that part

OpenStudy (texaschic101):

y - y1 = m(x - x1) slope(m) = -2/11 (-2,-1) x1 = -2 and y1 = -1 now we sub y - (-1) = -2/11(x - (-2) y + 1 = -2/11(x + 2) y + 1 = -2/11x - 4/11 y = -2/11x - 4/11 - 1 y = -2/11x - 4/11 - 11/11 y = -2/11x - 15/11

OpenStudy (anonymous):

apparently the awnser is y=-2/11x + 7/11

OpenStudy (anonymous):

thaats what my teacher wrote down

OpenStudy (texaschic101):

let me try again.... y - y1 = m(x - x1) (9,-3) x1 = 9 and y1 = -3 now we sub y - (-3) = -2/11(x - 9) y + 3 = -2/11x + 18/11 y = -2/11x + 18/11 - 3 y = -2/11x + 18/11 - 33/11 y = -2/11x - 15/11 hmmm.. I am getting the same answer by using the other point as well.

OpenStudy (texaschic101):

I am sorry...I lost track of time...I have to go now. Get somebody else to check this if you want to make sure